5.4.36 E Question Help The average math SAT score is 524 with a standard deviation of 114. A particular high school claims that its students have unusually high math SAT scores. A random sample of 50 students from this school was selected, and the mean math SAT score was 557 Is the high school justified in its claim? Explain because tho 2screithnt,naml z-score is v since it within the range of a usual event, namely within of the mean of the sample means Round to two decimal places as needed)
Answer #1

as here std error=std deviation/(n)1/2=114/(50)1/2=16.122

z score=(X-mean)/std error)=(557-524)/16.122=2.05

|No beacause the z score 2.05 is unusual since it does not falls within.......... namely within 2 standard deviation of the mean of te sample means